7. (10 points) Find the explicit solution to 1 94 + 6y + y = 0 with initial condition y(0) = and y(0) = -2
Answer #1

biven that gyl & Gy ty=0 yo) = 1 and 40) = -2 he gf y(x) +6d90 +960) = dra Jby@ + CD YOU + YUPO (90° + CD +1) 9100) Yos compO yo ! y (o) = f = (Citoje I = -2013 & Ge Y(x) = (6,+Gx) () yo = -2 4 = 1 (₂ + 0 x 1 (- / 3 ) Ź + C y(O) = -2 = & C x1 ---

If you have any doubt then please comment.